• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 08:07:45 on July17,2003

    Entry number 104226

    keyword=owl shift summary

    delta scan run summary

    0000:0030 waiting for right HRS to set to P0 = 2.221 GeV
    0040 noticed that septum was set for P0 = 2.093 GeV
    0041 started +4% scan @ Multi Foil
    0053 raster = OFF
    0110 moved to Single/Long Foil
    0131 raster = ON
    0150:0209 changing to P0 = 2.178 GeV
    0222 raster = OFF
    0236 moved to Multi Foil
    0253 raster = ON
    0305:0320 changing to P0 = 2.136 GeV
    0331 raster = OFF
    0344 moved to Single/Long Foil
    0404 raster = ON
    0420:0435 changing to P0 = 2.093 GeV
    0435 moved to Ref Cell, N2, 100psig, raster = ON, 2 microA
    0446 MCC called: QE Measurement, see elog 1160949 & 1160951
    0455 moved to Single/Long Foil
    0513 raster = OFF
    0528 moved to Multi Foil, raster = ON
    0541 raster = OFF
    0550:0603 changed to P0 = 2.051 GeV
    0613 raster = ON
    0626 moved to Single/Long Foil
    0630 paged JP/Rob for upcoming controlled access
    0649 raster = OFF
    0705 delta scan done
    0710 request controlled access
    0730 = 42.5%
    0800 installed left arm carbon target, removed sieve slit

    changing P0 to +4% (2.221GeV)